# TinySAM: Pushing the Envelope for Efficient Segment Anything Model

### 21 Dec 2023 v1

@@ Abstract
Recently segment anything model (SAM) has shown powerful segmentation capability and has drawn great attention in computer vision fields. Massive following works have developed various applications based on the pretrained SAM and achieved impressive performance on downstream vision tasks. However, SAM consists of heavy architectures and requires massive computational capacity, which hinders the further application of SAM on computation constrained edge devices. To this end, in this paper we propose a framework to obtain a tiny segment anything model (TinySAM) while maintaining the strong zero-shot performance. We first propose a full-stage knowledge distillation method with online hard prompt sampling strategy to distill a lightweight student model. We also adapt the post-training quantization to the promptable segmentation task and further reduce the computational cost. Moreover, a hierarchical segmenting everything strategy is proposed to accelerate the everything inference by 2×2$\times$2 × with almost no performance degradation. With all these proposed methods, our TinySAM leads to orders of magnitude computational reduction and pushes the envelope for efficient segment anything task. Extensive experiments on various zero-shot transfer tasks demonstrate the significantly advantageous performance of our TinySAM against counterpart methods. Pre-trained models and codes will be available at https://github.com/xinghaochen/TinySAM and https://gitee.com/mindspore/models/.
@@ 1 Introduction
Figure 1: (a) The overall framework of our proposed method. Consisting the modules of the full-stage knowledge distillation, the post training quantization and the hierarchical everything inference, the computation cost is down-scaled by magnitudes. (b) The proposed TinySAM can save considerable computation cost while maintaining the performance.
To further push the envelope for efficient segment anything model, in this paper we propose a full framework to obtain TinySAM that greatly reduces the computational cost while maintaining the zero-shot segmentation ability to maximum extent. Specifically, we propose a full-stage knowledge distillation method to improve the capability of the compact student network. The student network is distilled in an end-to-end manner with the supervision of teacher network from different network stages. Besides, a online hard prompt sampling strategy is proposed to make the distillation process attend more to hard examples and thus improves the final performance. We also adapt the post-training quantization to the promptable segmentation task and further reduce the computational cost. Moreover, we find that it takes tremendous computational cost for segmenting everything in an image since massive masks have to be generated from grid prompt points. To this end, a hierarchical segmenting everything strategy is proposed to accelerate the everything inference by 2×2$\times$2 × with almost no performance degradation. With all these proposed methods, our TinySAM leads to orders of magnitude computational reduction and pushes the envelope for efficient segment anything task. For example, TinySAM can achieve 100$\times$ acceleration for segment anything task compared with the original SAM. Extensive experiments on various zero-shot transfer tasks demonstrate the significantly advantageous performance of our TinySAM against counterparts.
@@ 3.1 Overview of TinySAM
This paper proposes a framework to get an highly efficient SAM, as described in Figure 1. Firstly, we introduce a full-stage knowledge distillation specifically designed for SAM in Section 3.2. To further activate the distillation process, an online hard prompt sampling is utilized to mine the hard knowledge from the teacher network to the student network. Secondly, a post-training quantization method described in Section 3.3 is adapted to promptable segmentation task and applied to the light-weight student network. Thirdly, a hierarchical everything inference mode described in 3.4 is designed for segmenting anything task, which can avoid massive redundant computation only with negligible accuracy loss and speedup the inference time by 2×2$\times$2 ×.
Figure 2: The framework of the full-stage knowledge distillation, for simplicity, prompt encoder is not shown here, which is frozen for the distillation process.
3.2 Full-Stage Knowledge Distillation
SAM consists of three sub-networks, i.e. image encoder, prompt encoder and mask decoder. The image encoder network is based on vision transformer [12] and consumes great computation cost. Inspired by MobileSAM [49], we use the lightweight TinyViT [46] to replace the original heavy image encoder network. Considerable performance decay exists for this simple substitution. Therefore, we propose a full stage knowledge distillation strategy to guide the lightweight image encoder during learning procedure from multiple knowledge level.
@@ 3.2 Full-Stage Knowledge Distillation
Generally, random sampling from labeled training data could be adopted to generate the prompts to drive the end-to-end training of prompt-based mask prediction network as SAM. To further ease the learning process of the distillation between teacher and lightweight student network, we propose a hard prompt sampling strategy, which makes the training samples concentrate in the difficult area for prediction. Taking points prompt as an example, points $P_{0}$ are initially sampled inside the labeled mask region $M_{gt}$. These initial points are fed into the network with input image to get the predicted mask region $M_{0}$. Then we sample the prompt points from the difference set of $M_{gt}$ and $M_{0}$, and we conduct the procedure interactively. The $(\mathit{i}+1)$-th round sampling points $P_{i}$ are sampled from the difference set of $M_{gt}$ and $M_{i}$, that is,
@@
When applied on the training process, the $i$-th iteration is random sampled from 00 to $9$, which makes the difficulty of sampled prompts in a constrained range.
@@ 3.4 Hierarchical Segmenting Everything
COCO [27] LVIS v1 [17]
Method FLOPs (G) AP APS APM APL AP APS APM APL
ViTDet-H [23] - 51.0 32.0 54.3 68.9 46.6 35.0 58.0 66.3
zero-shot transfer methods (segmentation module only):
SAM-H [22] 3166 46.5 30.8 51.0 61.7 44.7 32.5 57.6 65.5
SAM-L [22] 1681 45.5 30.2 50.1 60.5 43.5 31.1 56.3 65.1
SAM-B [22] 677 41.0 28.5 45.5 53.4 40.8 29.1 52.8 60.7
FastSAM [50] 344 37.9 23.9 43.4 50.0 34.5 24.6 46.2 50.8
MobileSAM [49] 232 41.0 24.4 44.5 58.6 37.0 24.7 47.8 59.1
TinySAM (Ours) 232 41.9 26.0 45.6 58.3 38.6 26.6 49.9 60.1
Q-TinySAM (Ours) 61 41.3 25.1 44.9 58.1 37.2 25.1 48.6 59.3
Table 1: Zero-shot instance segmentation results on COCO and LVIS v1 dataset. Zero-shot transfer methods are prompted with the detection boxes from fully-supervised ViTDet model. TinySAM and quantized Q-TinySAM demonstrate advantageous performance on average precision.
@@ 4.1 Implementation Details
We utilize the TinyViT-5M [46] as the lightweight student image encoder and SAM-H as the teacher model, following prior work [49]. Only $1\%$ of SA-1B dataset is used as the training data for full-stage distillation. We adopt Adam optimizer and train the student network for 8 epochs. For each iteration, we sample 64 prompts according to online hard prompt sampling strategy described in Section 3.2. To accelerate the distillation process, the image embeddings from the teacher network have been computed and stored in advance. Therefore, the heavy image encoder of teacher network is not necessary to compute repeatedly during training time. For post training quantization, we quantize all convolution layers, linear layers, deconvolution layers and matrix multiply layers. In convolution layers and deconvolution layers, we use channel-wise scaling factors. And for linear layers and matrix multiply layers, we apply layer-wise scaling factors and head-wise scaling factors respectively. For iterative search, we set $\alpha=0.01,\beta=1.2,n=100,rounds=3$. We calibrate quantized model on SA-1B dataset using 8 images. We conduct zero-shot evaluation on downstream tasks like instance segmentation and point prompt segmentation. Following the suggestions by SAM [21], the multi-output mode is adopted and the final mask prediction is the one with highest IoU prediction.
@@ 4.2 Zero-Shot Instance Segmentation
For zero-shot instance segmentation task, we strictly follow the experimental settings of SAM and use the object detection results of ViTDet-H [23] as the box prompt for instance segmentation. We evaluate the zero-shot instance segmentation task for models on the benchmark of COCO [28] dataset and LVIS v1 [17]. We compare our TinySAM with different variants of SAM [22], and also with prior efficient models like FastSAM [50] and MobileSAM [49]. As shown in Table 1, the proposed TinySAM obtained superior performance when compared with prior methods. Specifically, our TinySAM outperforms FastSAM [50] in terms of FLOPs and instance segmentation accuracy, i.e., about $4\%$ AP improvement with only $67\%$ FLOPs. With the same computational cost, our TinySAM also achieves $0.9\%+$ AP on COCO dataset than MobileSAM [49] and $1.6\%+$ AP on LVIS v1 dataset, respectively. Our W8A8 quantized variant of TinySAM (Q-TinySAM) also obtains competitive performance across different methods. Specifically, Q-TinySAM achieves $0.3\%+$ AP on COCO dataset than MobileSAM [49], with only $26\%$ FLOPs. Some visual results on COCO validation set are shown in Figure 5. Our proposed TinySAM captures more clear and smooth boundaries compared with other efficient variants of SAM.
@@ 4.3 Zero-shot Points Valid Mask Evaluation
Strategy Model mIoU (%) Time (s)
@@ 4.4 Everything Mode Acceleration
We evaluate our proposed hierarchical everything inference strategy on COCO validation set. All latency benchmarks are conducted on a single NVIDIA V100 GPU. We sample 100 images with the least img_id from val2017 and conduct everything mode inference on these samples. The threshold values used in the everything mode are all kept the same as default. The results are shown in Table 2. We apply the same threshold and stability score on the same model evaluated with different strategies to make a fair comparison. But they can be different between these models. Our hierarchical strategy achieves comparable results compared with original $32\times 32$ points grid strategy while the cost of inference time is reduced by about $50\%$. Figure 6 shows the visual comparison between points grid strategy and the proposed hierarchical strategy. We can find that the proposed hierarchical strategy avoids the over-finegrained segmentation of objects in some cases which makes up for some losses of IoU. More visual results of the efficient everything mode inference is shown in Figure 8.
Figure 7: Location change of the sampling prompts with iterations, the green star denote the new sampled point prompt with online hard prompt sampling strategy. With more iterations, the sampling points are more close to the edge region of the groundtruth mask
Figure 8: Visualization results of the proposed hierarchical everything inference on TinySAM model.
@@ 4.5 Ablation Studies
Impacts of different modules. We first evaluate the effects of different modules, i.e., full-stage knowledge distillation loss, online hard prompt sampling and post quantization, respectively. As shown in Table 3, utilizing our proposed full-stage distillation strategy improve the performance from $40.5\%$ to $41.4\%$. Further incorporated with the online hard prompt sampling strategy, our method could obtain another $0.5\%$ AP gain. Using post-training quantization results in $0.8\%$ AP degradation but also greatly reduces the computational cost, which is also discussed in the main experiments. As shown in Figure 7, we show the sampling process of our hard prompt sampling designed for full-stage knowledge distillation. The star symbols represent sampling point with different iterations. With some iterations, the sampling region is more closed to the edge of the target mask, which makes the prompt relatively harder for student networks to learn.
@@
2 + Online Hard Prompt Sampling 41.9
3 + Quantization 41.3
@@
Impacts of different distillation losses. For detailed full-stage knowledge distillation process, we investigate the necessity of the proposed three-level distillation from the teacher network. Table 4 shows the ablation results with different combinations of distillation losses. The output distillation loss takes important part since it is close to the supervision information and the similarity with teacher network directly reflects in the evaluation metric. Token loss and embedding loss both prove to be beneficial since they are related to key nodes of teacher network, which reflects the image-level information and the interaction of prompts with the image, respectively.
@@
Different bits for quantization. We here explore the influence of different bit width. The Table 5 reports the average precision on COCO dataset. From the results, we can conclude that quantization to $8$ bits results in only slight performance drop. We also demonstrate the performance by further reducing the quantization bit width to $6$ bits.
@@
MobileSAM 41.0 232
+ W8A8 39.8 61
+ W6A6 36.3 46
TinySAM (Ours) 41.9 232
+ W8A8 41.3 61
+ W6A6 38.6 46

### 09 Mar 2024 v2

@@ Abstract
Recently segment anything model (SAM) has shown powerful segmentation capability and has drawn great attention in computer vision fields. Massive following works have developed various applications based on the pretrained SAM and achieved impressive performance on downstream vision tasks. However, SAM consists of heavy architectures and requires massive computational capacity, which hinders the further application of SAM on computation constrained edge devices. To this end, in this paper we propose a framework to obtain a tiny segment anything model (TinySAM) while maintaining the strong zero-shot performance. We first propose a full-stage knowledge distillation method with hard prompt sampling and hard mask weighting strategy to distill a lightweight student model. We also adapt the post-training quantization to the promptable segmentation task and further reduce the computational cost. Moreover, a hierarchical segmenting everything strategy is proposed to accelerate the everything inference by 2×2$\times$2 × with almost no performance degradation. With all these proposed methods, our TinySAM leads to orders of magnitude computational reduction and pushes the envelope for efficient segment anything task. Extensive experiments on various zero-shot transfer tasks demonstrate the significantly advantageous performance of our TinySAM against counterpart methods. Pre-trained models and codes are available at https://github.com/xinghaochen/TinySAM and https://gitee.com/mindspore/models/tree/master/research/cv/TinySAM.
@@ 1 Introduction
Figure 1: (a) The overall framework of our proposed method. Consisting the modules of the hard mining full-stage knowledge distillation, the post training quantization and the hierarchical everything inference, the computation cost is down-scaled by magnitudes. (b) The proposed TinySAM can save considerable computation cost while maintaining the performance. The latency is tested with TensorRT on NVIDIA T4 GPU.
To further push the envelope for efficient segment anything model, in this paper we propose a full framework to obtain TinySAM that greatly reduces the computational cost while maintaining the zero-shot segmentation ability to maximum extent. Specifically, we propose a hard mining full-stage knowledge distillation method to improve the capability of the compact student network. The student network is distilled in an end-to-end manner with the supervision of teacher network from different network stages. A mask-weighted distillation loss is proposed to efficiently transfer the information from teacher to student through massive various SA-1B masks. Besides, an online hard prompt sampling strategy is proposed to make the distillation process attend more to hard examples and thus improves the final performance. We also adapt the post-training quantization to the promptable segmentation task and further reduce the computational cost. Moreover, we find that it takes tremendous computational cost for segmenting everything in an image since massive masks have to be generated from grid prompt points. To this end, a hierarchical segmenting everything strategy is proposed to accelerate the everything inference by 2×2$\times$2 × with almost no performance degradation. With all these proposed methods, our TinySAM leads to orders of magnitude computational reduction and pushes the envelope for efficient segment anything task. For example, TinySAM can achieve 100$\times$ acceleration for segment anything task compared with the original SAM. Extensive experiments on various zero-shot transfer tasks demonstrate the significantly advantageous performance of our TinySAM against counterparts.
@@ 2.1 Segment Anything Model
Figure 2: The framework of the hard mining full-stage knowledge distillation. For the massive masks of SA-1B dataset, we design the hard prompt sampling for prompts and hard mask weighting for distillation loss. For hard prompt sampling process, the stars represent sampling point with different iterations. With iterations of prompt sampling, the sampling region is more closed to the edge of the target mask, which makes the prompt relatively harder for student network to learn. Moreover, according to the gap between student and teacher network, different weight is assigned to each mask when calculating the distillation loss.
@@ 3.1 Overview of TinySAM
This paper proposes a framework to get an highly efficient SAM, as described in Figure 1. Firstly, we introduce a hard mining full-stage knowledge distillation specifically designed for SAM in Section 3.2. To further activate the distillation process, the proposed hard mask weighting and hard prompt sampling strategy are both utilized to mine the essential knowledge from the teacher network to the student network. Secondly, a post-training quantization method described in Section 3.3 is adapted to promptable segmentation task and applied to the light-weight student network. Thirdly, a hierarchical everything inference mode described in 3.4 is designed for segmenting anything task, which can avoid massive redundant computation only with negligible accuracy loss and speedup the inference time by 2×2$\times$2 ×.
3.2 Hard Mining Full-Stage Knowledge Distillation
SAM consists of three sub-networks, i.e. image encoder, prompt encoder and mask decoder. The image encoder network is based on vision transformer [12] and consumes great computation cost. Inspired by MobileSAM [50], we use the lightweight TinyViT [46] to replace the original heavy image encoder network. Considerable performance decay exists for this simple substitution. Therefore, we propose a hard mining full-stage knowledge distillation strategy to guide the lightweight image encoder during learning procedure from multiple knowledge levels.
@@ 3.2 Hard Mining Full-Stage Knowledge Distillation
Hard Mask Weighting. To make the knowledge distillation more effective, we design a hard mask weighting strategy when calculating the losses. There is an observation that masks could be extremely various in a single image of SA-1B dataset since the fine-grained granularity and no semantic constraints. As shown in Figure 2, segmenting the flag with complex boundary could be difficult while segmenting the rectangular window with high contrast color could be easy. The hard mask should reasonably be assigned with larger weight for student to learn. Specifically, we calculate the gap of student and teacher network output to indicate the mask hardness $\mathcal{H}_{i}$.
$\small\mathcal{H}_{i}=\mathrm{sigmoid}(\frac{\mathrm{IoU}({M}_{i}^{T},{M}_{i}^{GT})}{\mathrm{IoU}({M}_{i}^{S},{M}_{i}^{GT})+\epsilon}-1),$ (6)
where ${M}_{i}^{T}$, ${M}_{i}^{S}$, ${M}_{i}^{GT}$ represent the mask prediction of student network, the mask prediction of teacher network and the groundtruth for $i$th mask, respectively. Thus the distillation loss could be updated with
$\small\mathcal{L}_{distill}^{*}=\alpha*\mathcal{L}_{embedding}+\beta*\mathcal{L}_{token}+\gamma*\sum_{i=1}^{N}\mathcal{H}_{i}*\mathcal{L}_{output}^{i}.$ (7)
Hard Prompt Sampling. Generally, random sampling from labeled training data could be adopted to generate the prompts to drive the end-to-end training of prompt-based mask prediction network as SAM. To further ease the learning process of the distillation between teacher and lightweight student network, we propose a hard prompt sampling strategy, which makes the training samples concentrate in the difficult area for prediction. Taking points prompt as an example, points $P_{0}$ are initially sampled inside the labeled mask region $M_{gt}$. These initial points are fed into the network with input image to get the predicted mask region $M_{0}$. Then we sample the prompt points from the difference set of $M_{gt}$ and $M_{0}$, and we conduct the procedure interactively. The $(\mathit{i}+1)$-th round sampling points $P_{i}$ are sampled from the difference set of $M_{gt}$ and $M_{i}$, i.e.
@@
When applied on the training process, the $i$-th iteration is random sampled from 00 to $9$, which makes the difficulty of sampled prompts in a constrained range. The bottom of Figure 2 shows the location change of the sampling prompts with iterations, the green stars denote the new sampled point prompt with online hard prompt sampling strategy. With more iterations, the sampling points are more close to the edge region of the groundtruth mask.
@@ 3.3 Quantization
COCO [27] LVIS v1 [17]
Method FLOPs Lat.(ms) AP APS APM APL AP APS APM APL
ViTDet-H [23] - - 51.0 32.0 54.3 68.9 46.6 35.0 58.0 66.3
zero-shot transfer methods (segmentation module only):
SAM-H [22] 2976G 2392 46.6 30.8 51.0 61.7 44.7 32.5 57.6 65.5
SAM-L [22] 1491G 1146 46.2 30.2 50.1 60.5 43.5 31.1 56.3 65.1
SAM-B [22] 487G 368.8 43.4 28.5 45.5 53.4 40.8 29.1 52.8 60.7
FastSAM [51] 344G 153.6 37.9 23.9 43.4 50.0 34.5 24.6 46.2 50.8
EfficientSAM-Ti [47] 106G 81.0 42.3 26.7 46.2 57.4 39.9 28.9 51.0 59.9
MobileSAM [50] 42.0G 38.4 41.0 24.4 44.5 58.6 37.0 24.7 47.8 59.1
TinySAM (Ours) 42.0G 38.4 42.3 26.3 45.8 58.8 38.9 27.0 50.3 60.2
Q-TinySAM (Ours) 20.3G 24.0 41.4 25.6 45.1 57.9 38.5 26.6 49.8 59.8
Table 1: Zero-shot instance segmentation results on COCO and LVIS v1 dataset. Zero-shot transfer methods are prompted with the detection boxes from fully-supervised ViTDet model. TinySAM and quantized Q-TinySAM demonstrate advantageous performance on average precision. The latency is tested on NVIDIA T4 GPU.
@@ 4.1 Implementation Details
We utilize the TinyViT-5M [46] as the lightweight student image encoder and SAM-H as the teacher model, following prior work [50]. Partial SA-1B dataset is used as the training data for full-stage distillation. We adopt Adam optimizer and train the student network for 8 epochs. For each iteration, we sample 64 prompts according to hard prompt sampling strategy described in Section 3.2. To accelerate the distillation process, the image embeddings from the teacher network have been computed and stored in advance. Therefore, the heavy image encoder of teacher network is not necessary to compute repeatedly during training time. For post training quantization, we set $\alpha=0.01,\beta=1.2,n=100,rounds=3$ for iterative search. We calibrate quantized model on SA-1B dataset using 8 images. We conduct zero-shot evaluation on downstream tasks like instance segmentation and point prompt segmentation. Following the suggestions by SAM [21], the multi-output mode is adopted and the final mask prediction is the one with highest IoU prediction.
@@ 4.2 Zero-Shot Instance Segmentation
For zero-shot instance segmentation task, we strictly follow the experimental settings of SAM and use the object detection results of ViTDet-H [23] as the box prompt for instance segmentation. We evaluate the zero-shot instance segmentation task for models on the benchmark of COCO [28] dataset and LVIS v1 [17]. We compare our TinySAM with different variants of SAM [22], and also with prior efficient models like FastSAM [51], MobileSAM [50] and EfficientSAM [47]. As shown in Table 1, the proposed TinySAM obtained superior performance when compared with prior methods. Specifically, our TinySAM outperforms FastSAM [51] in terms of FLOPs and instance segmentation accuracy, i.e., about $4\%$ AP improvement with only $12.2\%$ FLOPs and $25\%$ latency. With the same computational cost, our TinySAM also achieves $1.3\%+$ AP on COCO dataset than MobileSAM [50] and $1.9\%+$ AP on LVIS v1 dataset, respectively. With similar performance on COCO dataset, TinySAM is 2×2$\times$2 × faster than EfficientSAM [47]. Our W8A8 quantized variant of TinySAM (Q-TinySAM) also obtains competitive performance across different methods. Specifically, Q-TinySAM achieves $0.4\%+$ AP on COCO dataset than MobileSAM [50], with only $48\%$ FLOPs and $62.5\%$ latency. Some visual results on COCO validation set are shown in Figure 5. Our proposed TinySAM captures more clear and smooth boundaries compared with other efficient variants of SAM.
@@ 4.3 Zero-shot Points Valid Mask Evaluation
Strategy Model mIoU Time (s)
@@ 4.4 Everything Mode Acceleration
We evaluate our proposed hierarchical everything inference strategy on COCO validation set. Latency benchmarks are conducted on a single NVIDIA V100 GPU for everything mode. We sample 100 images with the least img_id from val2017 and conduct everything mode inference on these samples. The threshold values used in the everything mode are all kept the same as default. The results are shown in Table 2. We apply the same threshold and stability score on the same model evaluated with different strategies to make a fair comparison. But they can be different between these models. Our hierarchical strategy achieves comparable results compared with original $32\times 32$ points grid strategy while the cost of inference time is reduced by about $50\%$. Figure 6 shows the visual comparison between points grid strategy and the proposed hierarchical strategy. We can find that the proposed hierarchical strategy avoids the over-finegrained segmentation of objects in some cases which makes up for some losses of IoU. More visual results of the efficient everything mode inference is shown in Figure 7.
Figure 7: Visualization results of hierarchical everything inference with TinySAM model.
@@ 4.5 Ablation Studies
Impacts of different modules. We first evaluate the effects of different modules, i.e., full-stage knowledge distillation loss, online hard prompt sampling and post quantization, respectively. As shown in Table 3, utilizing our proposed full-stage distillation strategy improve the performance from $40.7\%$ to $41.4\%$. Incorporated with the online hard prompt sampling strategy, our method could obtain $0.5\%$ AP gain. With the hard mask weighting loss, the performance can further increase to $42.3\%$. Using post-training quantization results in $0.9\%$ AP degradation but also greatly reduces the computational cost, which is also discussed in the main experiments.
@@
2 + Hard Prompt Sampling 41.9
3 + Hard Mask Weighting 42.3
4 + Quantization 41.4
@@
Impacts of different distillation losses. For detailed full-stage knowledge distillation process, we investigate the necessity of the proposed three-level distillation from the teacher network. Table 4 shows the ablation results with different combinations of distillation losses. The output distillation loss takes important part since it is close to the supervision information and the similarity with teacher network directly reflects in the evaluation metric. Token loss and embedding loss both prove to be beneficial since they are related to key nodes of teacher network, which reflects the image-level information and the interaction of prompts with the image, respectively. Hard mask weighting for output loss can further boost the performance.
@@
Point density and threshold for hierarchical strategy. In Table. 5, we conduct ablation study with different settings of point density and high-confidence mask threshold $\tau$. More points and higher threshold $\tau$ lead to more precise results but longer inference time. The point density of 2nd round is more sensitive compared to the 1st one. Considering both accuracy and efficiency, the setting in bold is a good balance and used for other experiments of everything inference mode.
Points per side 1st/2nd Thresh. $\tau$ mIoU Time (s)
4/16 8.5 0.5521 0.3571
8/32 8.5 0.6061 0.9303
10/32 8.5 0.6078 1.2774
8/32 7.0 0.6018 0.8154
8/32 10.0 0.6067 1.1819
32/- 0.6137 1.7790
Table 5: Ablation on point density and threshold for hierarchical strategy.
Different bits for quantization. We here explore the influence of different bit width. Table 6 reports the average precision on COCO dataset. From the results, we can conclude that quantization to $8$ bits results in only slight performance drop. We also demonstrate the performance by further reducing the quantization bit width to $6$ bits.
@@
MobileSAM 41.0 42.0
+ W8A8 39.8 20.28
+ W6A6 36.3 18.45
TinySAM (Ours) 42.3 42.0
+ W8A8 41.4 20.28
+ W6A6 39.0 18.45